# BannerAgency: Advertising Banner Design with Multimodal LLM Agents

### v1

@@ 5.2 Results and Analysis
We present quantitative and visual comparisons in  Tab. 1(a) and  Fig. 8, respectively. BannerAgency outperforms all baselines across all aspects. Baseline limitations include: fixed-size generation preventing non-standard dimensions like $300\times 250$, failure to incorporate provided logos, and incorrect text rendering requiring correction [58]. Among BannerAgency variants, our preference heatmap in  Tab. 1(b) shows Claude3.5-Sonnet surpasses GPT-4o ($\mathbf{[D]}$ vs. $\mathbf{[E]}$ and $\mathbf{[B]}$ vs. $\mathbf{[C]}$), especially in SVG implementation. Human evaluators showed no preference between SVG and Figma implementations ($\mathbf{[A]}$ vs. $\mathbf{[B]}$). Notably, removing the Background Designer component (i.e., w/o BG) significantly impairs visual appeal, resulting in plain and unengaging designs, particularly with the GPT-4o backbone. This underscores the critical role of specialized background generation in creating visually compelling banner advertisements.
@@
Fig. 9(a) demonstrates the design improvement after the iterative refinement process. The initial designs are upgraded by enhancing content clarity, layout composition, and visual styling without radical redesigns. As shown in  Fig. 9(b), our evaluation reveals a clear upward trend in quality, with improvement scores steadily increasing across iterations and later versions receiving more votes. Notably, the vote distribution indicates that refinement effectiveness is design-context dependent—while the final iteration was most frequently preferred, many participants favored earlier iterations or the initial design. This finding highlights the value of maintaining a complete version track, allowing advertisers to select the specific iteration that best aligns with their vision, regardless of where it falls in the refinement sequence. This process also reflects how our component-based approach enables practical, editable design alternatives.
@@
After presenting versatility to various sizes in Fig. 1, we showcase in  Fig. 10 two more examples highlighting cross-genre aesthetic adaptation and cross-cultural linguistic flexibility.  Fig. 10(a) presents six aesthetic styles for the same automation company request. We observe how the design adapt coherently to each style while maintaining color schemes that complement the AutoMate logo. In  Fig. 10(b), we show how the system successfully maintains the 20% discount promotion across all versions while accommodating language-specific conventions and cultural preferences, which directly addresses the industry challenge of creating culturally appropriate marketing materials without specialized design teams. Empowered by the world knowledge of multimodal LLMs, our training-free BannerAgency demonstrates remarkable adaptability to diverse user requirements without any additional fine-tuning or domain-specific datasets.
6 Conclusion and Future Work
@@
We introduced BannerAgency, a training-free MLLM agent system that transforms advertising banner design by simulating professional design teams through specialized agents (Strategist, Background Designer, Foreground Designer, and Developer). Our approach generates component-based, editable designs in industry-standard formats from multimodal inputs without domain-specific fine-tuning, overcoming the limitations of pixel-based methods. Through our BannerRequest400 benchmark featuring 100 unique logos paired with 400 diverse banner requests, we demonstrate BannerAgency’s effectiveness in producing high-quality, versatile, and editable designs that kickstart the creative design process. This represents a significant step toward democratizing professional-grade design capabilities across the digital advertising landscape. In the future, we aim to extend BannerAgency by: (1) expanding multimodal input to incorporate product images, (2) introducing more sophisticated decorative elements beyond basic geometric shapes, 3) allowing for more hierarchical architecture for better collaboration among agents, and (4) considering conversion metric optimization to directly connect design choices with advertising performance.
Supplementary Material
@@
In this appendix, we present more details on our experimental setup ( Sec. 7), human study setup ( Sec. 8), text-free background generation ( Sec. 9), and foreground variation generation ( Sec. 9). We also show detailed statistics results on the refinement human study ( Sec. 10) and general graphic design evaluation ( Sec. 10). The detailed prompt for text-to-image generation models in Fig. 8 of our main paper is presented in  Sec. 11. Lastly, we showcase more results from our BannerAgency in  Sec. 12.
7 More Details on Experimental Setup
@@
We present a detailed cost breakdown in USD for generating banner advertisements of size $300\times 250$ pixels in  Tab. 3. Our BannerAgency upscales images to $1024\times 864$ pixels to maintain visual quality before rendering. Costs are calculated based on Claude 3.5 Sonnet API pricing (\$3/MTok for input, \$15/MTok for output) and FLUX-1 [schnell] image generation (\$0.0027/MP). We compare both Figma and SVG implementation approaches across three scenarios: single image generation, generation with four-rounds of iterative refinement, and multi-variation generation with four design alternatives as detailed in  Sec. 9. The dramatic difference between the two implementations stems from fundamental implementation constraints: while SVG can be generated directly as code that LLMs frequently encounter in training data, Figma plugin JavaScript is uncommon in LLM training corpora. To ensure reliable execution, our Figma implementation requires providing templates where the Developer agent only needs to call predefined functions, significantly increasing token usage. Notably, both implementation approaches offer extraordinary cost-effectiveness compared to hiring professional designers, who typically charge \$50-150 per hour [51]. Even our most expensive scenario (Figma with feedback at \$1.213) represents less than 2.5% of traditional design costs, offering transformative economics for advertising campaigns requiring multiple banner variations.
@@
create_unique_image_name Generates a unique filename by combining a timestamp with the banner name to ensure distinct identification of each banner design output (e.g., ”20240328_154227_kids_dentist_clinic_banner”).
@@
8 More Details on Human Study
@@
9 More Details on Design Generation
@@
Algorithm 1 shows the detailed behavior of the text-free background generation.
@@
Input: User requirements $R$, Logo characteristics $L$
Output: Background image path $P$, Description $D$ (if generated)
Agent Tools:
$\textsc{FindImagePath}:R\mapsto(P,\text{bool})$Primary search tool
[…]
@@
1:  $P,path\_exists\leftarrow\textsc{FindImagePath}(R)$
2:  if $path\_exists$ then
@@
7:     while $contains\_text$ AND $attempts<5$ do
@@
12:        if $contains\_text$ then
@@
10 More Results Analysis
@@
To complement Fig.9(a) in our main paper, we present the complete design evolution in  Fig. 14. We also present the full statistics from human study on the refinement process in  Tab. 6. We observe a statistically significant improvement in quality ratings, increasing from $2.56$ ($\sigma=0.89$) for initial designs to $3.55$ ($\sigma=0.77$) for refined versions ($p<0.001$).
@@
Table 6: Full statistics of refinement human study
@@
Table 7: Evaluation metrics for generic design quality. Primary metrics include: Design and Layout (DL), Content Relevance (CR), Typography and Color (TC), Graphics and Images (GI), and Innovation (INV) [31], as well as technical aspects [23] such as Alignment (ALM), Overlap prevention (OLP), and White Space utilization (WS). All metrics are evaluated on a 10-point scale by GPT-4o, except for FID (Fréchet Inception Distance), which compares the similarity between generated designs and exemplar ones.
@@
We compute general graphic design metrics [31, 23] in  Tab. 7 to provide complementary analysis. Following PPTAgent [75], we compute the FID (Fréchet Inception Distance) score between the generated banner ad designs (400) and exemplar designs (3355)https://huggingface.co/datasets/PeterBrendan/AdImageNet using 64-dimensional feature vectors. We observe that Figma implementations generally achieve lower FID scores (GPT-4o: 2.60, Claude3.5-Sonnet: 4.99) than SVG implementations, suggesting closer visual similarity to real-world examples. The evaluation metrics, all adopted from prior work, include GPT-4o assessment of five primary aspects: (i) design and layout (DL), (ii) content relevance (CR), (iii) typography and color (TC), (iv) graphics and images (GI), and (v) innovation (INV)  [31], alongside technical metrics on alignment (ALM), overlap prevention (OLP), and white space utilization (WS) [23]. The results reveal that BannerAgency implementations are competitive with or surpass baseline methods across most metrics, particularly in content relevance (CR) where Claude3.5-Sonnet achieves 8.43, exceeding all baselines. The SVG implementation with Claude3.5-Sonnet notably outperforms in white space utilization (5.51) compared to baselines (highest: 4.53 with OpenCOLE). These results are presented in the supplementary materials as these generic design metrics, while informative, don’t fully capture banner advertisement-specific requirements such as attention guidance, call-to-action effectiveness, and brand consistency, which are more central to our paper’s focus.
11 Detailed Prompt in Fig.8 of Main Paper
@@
    “Please create a 300x250 banner ad image with a background that features a subtle, sophisticated gradient of light blue and white, evoking a sense of calm and intellectual engagement. Overlay this with a faint, abstract pattern of interconnected nodes and lines, symbolizing the complex network of AI and ethical considerations. Centered at the top, include the text ‘Join the Discussion on AI Ethics’ in a bold, dark gray sans-serif font. Below this, in a slightly smaller font, add ‘Share Your Expert Insights’ in the same blue color as our logo. At the bottom center, place a prominent CTA button in blue [#4285F4] with white text that reads ‘Contribute Now’. Position our logo in the bottom left corner, ensuring it is clearly visible but not overpowering the main message.”
@@
    “Create a 300x250 banner ad image with a background showcasing a vibrant, high-quality selection of premium cuts of meat, professionally prepared and arranged to look fresh and enticing. Use warm, inviting lighting to emphasize the quality and freshness of the products. Overlay the text ‘20% Seasonal Discount on Premium Cuts’ in bold, white font at the top center of the banner. Below this, include the text ‘Shop Local, Eat Fresh’ in a slightly smaller, italicized white font. Place a prominent, eye-catching CTA button at the bottom center of the banner with the text ‘Visit Us Today’ in bold, dark red [#C8102E] font on a white background. Position the Blade & Block logo in the bottom left corner, ensuring it is clearly visible but not overpowering the main message.”
  3. 3.
    “Create a 300x250 banner ad image with a romantic Valentine’s Day theme. The background should feature a soft, dreamy blend of pink and red hues with subtle heart patterns. In the center, place an image of a beautifully wrapped gift box with a ribbon, symbolizing unique gifts. Overlay the text ‘Happy Valentine’s Day!’ in an elegant, large serif font at the top. Below that, in a slightly smaller but bold font, add ‘Special Offer for the First 50 Customers!’ and ‘Unique Gifts for Your Loved One’ to highlight the promotion. Place a prominent CTA button at the bottom center with the text ‘Shop Now’ in white on a dark brown [#4A2E2B] background. Position the Love & Wrapped logo in the bottom right corner, ensuring it is clearly visible but not overpowering the main message.”
  4. 4.
@@
    “Please create a 300x250 banner ad image for a dating app targeting single adults aged 18-35. Use a bright and eye-catching color scheme with a background image that features a vibrant, romantic scene such as a sunset beach or a cityscape with couples walking hand-in-hand. Overlay the text ‘Swipe into Romance’ prominently in a bold, playful font, ensuring it stands out. Below this, include a CTA button with the text ‘Download Now’ in a contrasting color to grab attention. Place the Swoon logo in the bottom right corner, ensuring it is clearly visible but does not overpower the main message. The overall design should be modern, clean, and appealing to the target audience, encouraging them to download the app.”
12 More Visual Results

### v2

@@ 5.2 Results and Analysis
We present quantitative and visual comparisons in Tab. 5.1 (a) and Fig. 7, respectively. BannerAgency outperforms all baselines across all aspects. Baseline limitations include: fixed-size generation preventing non-standard dimensions like $300\times 250$, failure to incorporate provided logos, and incorrect text rendering requiring correction Shimoda et al. (2024). Among BannerAgency variants, our preference heatmap in Tab. 5.1 (b) shows Claude3.5-Sonnet surpasses GPT-4o ($\mathbf{[D]}$ vs. $\mathbf{[E]}$ and $\mathbf{[B]}$ vs. $\mathbf{[C]}$), especially in SVG implementation. Human evaluators showed no preference between SVG and Figma implementations ($\mathbf{[A]}$ vs. $\mathbf{[B]}$). Notably, removing the Background Designer component (ie, w/o BG) significantly impairs visual appeal, resulting in plain and unengaging designs, particularly with the GPT-4o backbone. This underscores the critical role of specialized background generation in creating visually compelling banner advertisements.
@@
Fig. 8 (a) demonstrates the design improvement after the iterative refinement process. The initial designs are upgraded by enhancing content clarity, layout composition, and visual styling without radical redesigns. As shown in Fig. 8 (b), our evaluation reveals a clear upward trend in quality, with improvement scores steadily increasing across iterations and later versions receiving more votes. Notably, the vote distribution indicates that refinement effectiveness is design-context dependent. While the final iteration was most frequently preferred, many participants favored earlier iterations or the initial design. This finding highlights the value of maintaining a complete version track, allowing advertisers to select the specific iteration that best aligns with their vision, regardless of where it falls in the refinement sequence. This process also reflects how our component-based approach enables practical, editable design alternatives. The full refinement process is visualized in Appendix A.4.
@@
We showcase in Fig. 9 two more examples highlighting cross-genre aesthetic adaptation and cross-cultural linguistic flexibility, respectively. We observe how the designs adapt coherently to each style while maintaining color schemes that complement the AutoMate logo. For different linguistic requests with same canvas input, we show how the system successfully accommodates language-specific conventions and cultural preferences with tailored fonts and styles, which directly addresses the industry challenge of creating culturally appropriate marketing materials without specialized design teams. Empowered by the world knowledge of multimodal LLMs, our training-free BannerAgency demonstrates remarkable adaptability to diverse user requirements without any additional fine-tuning or domain-specific datasets.
6 Conclusion
@@
We introduced BannerAgency, a training-free MLLM agent system that transforms advertising banner design by simulating professional design teams through specialized agents (Strategist, Background Designer, Foreground Designer, and Developer). Our approach generates component-based, editable designs in industry-standard formats from multimodal inputs without domain-specific fine-tuning, overcoming the limitations of pixel-based methods. Through our BannerRequest400 benchmark featuring 100 unique logos paired with 400 diverse banner requests, we demonstrate BannerAgency’s effectiveness in producing high-quality, versatile, and editable designs that kickstart the creative design process. This represents a significant step toward democratizing professional-grade design capabilities across the digital advertising landscape.
Limitations
While our approach represents a significant step toward leveraging MLLM agents for versatile and editable graphic design generation, there exist some limitations. First, we exclusively experiment with logo images as multimodal input. It is sufficient for abstract service banner ads but neglects product images which are also crucial in advertisements. Second, the decorative elements are limited to simple geometric shapes due to the lack of finetuning or domain-specific guidance. Similarly, we do not address visual texts with artistic effect. These limitations reflect our primary objective: to validate design knowledge already embedded in MLLMs and demonstrate the agent system’s versatility across varying requests (banner sizes, styles, etc.) while maintaining output editability. In the future, we will dedicate efforts to the upgrade of each agent by equipping them with more advanced tools or further decompose different agents into hierarchical teams. To reiterate, we aim to extend BannerAgency by: (1) expanding multimodal input to incorporate product images, (2) introducing more sophisticated decorative elements beyond basic geometric shapes, (3) allowing for more hierarchical architecture for better collaboration among agents, and (4) considering conversion metric optimization to directly connect design choices with advertising performance.
Ethics Statement
The BannerAgency framework, while advancing automated banner design through MLLM agents that simulate professional design workflows, acknowledges important ethical considerations. We recognize potential concerns regarding employment impacts on human designers, conscious content generation that avoids harmful stereotypes or misleading advertising, intellectual property considerations when incorporating brand assets, transparency with end users about AI-generated content, and potential biases in design choices that may perpetuate inequities in representation. Our component-based approach deliberately preserves human creative oversight through editable outputs, ensuring designers maintain agency in the final creative process rather than being replaced. We commit to ongoing evaluation of these ethical dimensions, including regular bias assessments of generated content, and encourage the research community to further explore responsible deployment strategies for automated design systems that complement rather than supplant human creativity.
Appendix A Appendix
@@
In this appendix, we present more details on our experimental setup (Sec. A.1), human study setup (Sec. A.2), text-free background generation (Sec. A.3), and foreground variation generation (Sec. A.3). We also show detailed statistics results on the refinement human study (Sec. A.4) and general graphic design evaluation (Sec. A.4). The detailed prompt for text-to-image generation models in Fig. 7 (a) of our main paper is presented in Sec. A.5. The agent profiles are displayed in Sec. A.6 Lastly, we showcase more results from our BannerAgency in Sec. A.7.
A.1 More Details on Experimental Setup
@@
We present a detailed cost breakdown in USD for generating banner advertisements of size $300\times 250$ pixels in Tab. 3. Our BannerAgency upscales images to $1024\times 864$ pixels to maintain visual quality before rendering. Costs are calculated based on Claude 3.5 Sonnet API pricing (\$3/MTok for input, \$15/MTok for output) and FLUX-1 [schnell] image generation (\$0.0027/MP). We compare both Figma and SVG implementation approaches across three scenarios: single image generation, generation with four-rounds of iterative refinement, and multi-variation generation with four design alternatives as detailed in Sec. A.3. The dramatic difference between the two implementations stems from fundamental implementation constraints: while SVG can be generated directly as code that LLMs frequently encounter in training data, Figma plugin JavaScript is uncommon in LLM training corpora. To ensure reliable execution, our Figma implementation requires providing templates where the Developer agent only needs to call predefined functions, significantly increasing token usage. Notably, both implementation approaches offer extraordinary cost-effectiveness compared to hiring professional designers, who typically charge \$50-150 per hour ManyPixels (2025). Even our most expensive scenario (Figma with feedback at \$1.213) represents less than 2.5% of traditional design costs, offering transformative economics for advertising campaigns requiring multiple banner variations.
@@
Agent profile definition.
Please refer to Sec. A.6 for each agent’s profile definition.
@@
create_unique_image_name Generates a unique filename by combining a timestamp with the banner name to ensure distinct identification of each banner design output (e.g., "20240328_154227_kids_dentist_clinic_banner").
@@
A.2 More Details on Human Study
@@
A.3 More Details on Design Generation
@@
Alg. A.3 shows the detailed behavior of the text-free background generation.
@@
Input: User requirements $R$, Logo characteristics
Output: Background image path $P$, Description $D$ (if generated) Agent Tools: $\textsc{FindImagePath}:R\mapsto(P,\text{bool})$Primary search tool
$\textsc{T2I}:D\mapsto P$
@@
[…]
2:if path_existspath\_exists then
@@
7:  while contains_textcontains\_text AND $attempts<5$ do
@@
12:   if contains_textcontains\_text then
@@
A.4 More Results Analysis
Quantitative analysis of spatial errors in generated designs.
We provide quantitative analysis on canvas overflow - the most critical spatial error in banner design. Canvas overflow occurs when elements extend beyond canvas boundaries, rendering content unusable and causing production-blocking errors. We investigate on how frequently the canvas overflow problem happens in our generated design with GPT-4o and Claude-3.5 serving the agent backbone, respectively. The results are summarized in Tab. 6. These results demonstrate that BannerAgency maintains proper canvas boundaries for over 98% of generated elements, with Claude-3.5 Sonnet achieving particularly strong performance at 98.96% accuracy. This low overflow rate validates our system’s spatial constraint handling for the most critical requirement in professional banner design. We also want to emphasize that, thanks to our flexible editability of foreground elements, spatial errors could be easily fixed by just dragging the elements in a design interface such as Figma.
Metric GPT-4o Claude-3.5
# Designs 416 408
# Elements 2220 2510
# Overflow Elements 41 26
Overflow Rate (%) 1.85% 1.04%
Table 6: Canvas overflow analysis.
@@ A.4 More Results Analysis
To complement Fig. 8 (a) in our main paper, we present the complete design evolution in Fig. 13. We also present the full statistics from human study on the refinement process in Fig. 7. We observe a statistically significant improvement in quality ratings, increasing from $2.56$ ($\sigma=0.89$) for initial designs to $3.55$ ($\sigma=0.77$) for refined versions ($p<0.001$).
@@
Table 7: Full statistics of refinement human study.
@@
Table 8: Evaluation metrics for generic design quality. Primary metrics include: Design and Layout (DL), Content Relevance (CR), Typography and Color (TC), Graphics and Images (GI), and Innovation (INV) Inoue et al. (2024), as well as technical aspects Haraguchi et al. (2024) such as Alignment (ALM), Overlap prevention (OLP), and White Space utilization (WS). All metrics are evaluated on a 10-point scale by GPT-4o, except for FID (Fréchet Inception Distance), which compares the similarity between generated designs and exemplar ones.
@@
We compute general graphic design metrics Inoue et al. (2024); Haraguchi et al. (2024) in Tab. A.4 to provide complementary analysis. Following PPTAgent Zheng et al. (2025), we compute the FID (Fréchet Inception Distance) score between the generated banner ad designs (400) and exemplar designs (3355)https://huggingface.co/datasets/PeterBrendan/AdImageNet using 64-dimensional feature vectors. We observe that Figma implementations generally achieve lower FID scores (GPT-4o: 2.60, Claude3.5-Sonnet: 4.99) than SVG implementations, suggesting closer visual similarity to real-world examples. The evaluation metrics, all adopted from prior work, include GPT-4o assessment of five primary aspects: (i) design and layout (DL), (ii) content relevance (CR), (iii) typography and color (TC), (iv) graphics and images (GI), and (v) innovation (INV)  Inoue et al. (2024), alongside technical metrics on alignment (ALM), overlap prevention (OLP), and white space utilization (WS) Haraguchi et al. (2024). The results reveal that BannerAgency implementations are competitive with or surpass baseline methods across most metrics, particularly in content relevance (CR) where Claude3.5-Sonnet achieves 8.43, exceeding all baselines. The SVG implementation with Claude3.5-Sonnet notably outperforms in white space utilization (5.51) compared to baselines (highest: 4.53 with OpenCOLE). These results are presented in the supplementary materials as these generic design metrics, while informative, don’t fully capture banner advertisement-specific requirements such as attention guidance, call-to-action effectiveness, and brand consistency, which are more central to our paper’s focus.
A.5 Detailed Prompt in Fig.7 (a) of Main Paper
@@
    "Create a 300x250 banner ad image with a romantic Valentine’s Day theme. The background should feature a soft, dreamy blend of pink and red hues with subtle heart patterns. In the center, place an image of a beautifully wrapped gift box with a ribbon, symbolizing unique gifts. Overlay the text ‘Happy Valentine’s Day!’ in an elegant, large serif font at the top. Below that, in a slightly smaller but bold font, add ‘Special Offer for the First 50 Customers!’ and ‘Unique Gifts for Your Loved One’ to highlight the promotion. Place a prominent CTA button at the bottom center with the text ‘Shop Now’ in white on a dark brown [#4A2E2B] background. Position the Love & Wrapped logo in the bottom right corner, ensuring it is clearly visible but not overpowering the main message.”
@@
    "Please create a 300x250 banner ad image for a dating app targeting single adults aged 18-35. Use a bright and eye-catching color scheme with a background image that features a vibrant, romantic scene such as a sunset beach or a cityscape with couples walking hand-in-hand. Overlay the text ‘Swipe into Romance’ prominently in a bold, playful font, ensuring it stands out. Below this, include a CTA button with the text ‘Download Now’ in a contrasting color to grab attention. Place the Swoon logo in the bottom right corner, ensuring it is clearly visible but does not overpower the main message. The overall design should be modern, clean, and appealing to the target audience, encouraging them to download the app.”
A.6 Agent Prompts
Strategist You are an expert banner objective setter. When a user requests a banner, develop high-level objectives for an effective banner that would: - Support the banner’s primary purpose - Create appropriate mood and tone - Appeal to the target audience
Core Decision Flow: 1. First, check if a background image path is provided in the user requirement using FindImagePath tool  - If yes: Validate the path exists and return it  - If no: Proceed with background image generation For Background Image Generation (when no path provided): 1. Analyze the provided objectives and user requirements 2. Create a detailed background description considering:  - Visual style and mood  - Color schemes  - Composition elements  - Focal points  You should  1. Focus only on the background’s visual and structural elements, including composition, key visual areas, and how these areas interact with one another.  2. Avoid saying something like "reserved for a CTA button", "support text readability", "for a banner" which could lead text existence in the generated image. The background needs free of text elements.  3. Avoid generic or abstract language like "guiding visual focus" or "framing the foreground." These terms do not add valuable visual detail to the background description. 3. Generate the background using T2I tool with the description. To use this tool, you have to create unique image name for the background with create_unique_image_name tool. 4. State where the generated image is saved. Use the path returned from the T2I tool. 5. Evaluate if the generated image contain any texts using the TextChecker. If yes, repeat step 3-5 by regenerating the background until it reports no. Important Guidelines: - Focus exclusively on background design - Avoid including any text elements in the generated background - If no specific background requirements are provided, use the objectives and user requirements to determine an appropriate background design. Output: - For existing images: Provide validated image path - For new images: Provide the final image path as well as the exact description used for T2I tool
User requirements: {user_input} Banner Objectives: - Primary Purpose: {purpose} - Target Audience: {audience} - Mood and Tone: {mood}
Background Designer
LAYOUT PATTERNS AND IMPLEMENTATION: 1. LEFT/RIGHT CONTENT  - Left-Content Pattern:   * All text elements aligned 20-30% from left edge   * Headline: Top left quadrant   * Supporting text: Below headline   * CTA: Bottom of text stack   * Use left-aligned text  - Right-Content Pattern:   * All text elements aligned 60-70% from left edge   * Mirror of left-content pattern   * Use right-aligned text   * Maintain right margin for readability 2. READING PATTERN BASED  - Z-Pattern:   * Headline: Top left   * Key visual/feature: Top right   * Supporting text: Bottom left diagonal   * CTA: Bottom right   * Follow natural eye movement  - F-Pattern:   * Headline: Top bar   * First key point: Upper left   * Second key point: Middle left   * CTA: Bottom, aligned with points   * Perfect for scannable content 3. CENTERED ARRANGEMENTS  - Centered Pattern:   * All elements center-aligned   * Headline: Upper third   * Supporting text: Middle   * CTA: Lower third   * Maintain balanced spacing  - Circular Pattern:   * Central focal point   * Text elements radiate outward   * CTA at bottom of circle   * Create visual movement 4. DESIGN PRINCIPLE BASED  - Golden-Ratio Layout:   * Key elements at golden ratio points (1.618:1)   * Start from largest text   * Spiral outward for hierarchy   * CTA at natural endpoint  - Rule-of-Thirds:   * Align elements to grid intersections   * Headline: Top third   * Supporting text: Middle third   * CTA: Bottom third intersection 5. DYNAMIC LAYOUTS  - Diagonal Pattern:   * Elements follow diagonal line   * Top left to bottom right (usual)   * Or bottom left to top right   * Create dynamic movement  - Asymmetrical:   * Intentionally unbalanced   * Heavy elements offset by white space   * Create tension and interest   * Maintain readability 6. HIERARCHICAL LAYOUTS  - Top-Down Pattern:   * Clear vertical progression   * Each element below previous   * Consistent left alignment   * Decreasing text sizes  - Pyramid Pattern:   * Largest element at top   * Progressively smaller elements   * CTA as focal point at bottom   * Clear visual hierarchy 7. STRUCTURED LAYOUTS  - Grid Pattern:   * Elements aligned to modular grid   * Consistent spacing   * Clear columns and rows   * Professional and organized IMPLEMENTATION GUIDELINES: 1. Typography Hierarchy:  - Headline: 32-48px  - Subheadline: 24-32px  - Body text: 16-20px  - CTA: 18-24px 2. Spacing Rules:  - Minimum 20px between elements  - Maintain 40px from banner edges  - Scale spacing with banner size  - Use consistent multiples 3. Color and Contrast:  - Maintain 4.5:1 contrast ratio  - Use consistent color scheme  - Consider background variation  - Ensure CTA stands out 4. Responsive Considerations:  - Scale text proportionally  - Maintain relative spacing  - Adjust for banner dimensions  - Preserve hierarchy 5. Creative and matching font:  - Choose the most matching font for the purpose  - Try to be creative in font selection Always document your layout rationale and ensure all positions are precisely specified in pixels or percentages. Below are the examples for major layout styles: {json.dumps(layout_demonstrations, indent=2)} Analyze this banner background image ({width}x{height}px) as required and take the following into consideration.
User requirements: {user_input} Banner Objectives: - Primary Purpose: {purpose} - Target Audience: {audience} - Mood and Tone: {mood}
Provide complete specifications following the schema exactly.
Foreground Designer
1. Update ui.html with background and logo image names 2. Follow the JavaScript template 3. Generate and save plugin code with layout implementation 4. Render the image in Figma desktop and provide the file path output by the render_and_save_image tool. e.g. ../rendered/xxx.png 5. Backup plugin Use these tools in sequence: 1. update_image_list - Modify ui.html 2. read_plugin_template - Get template code 3. save_plugin_code - Save code.js 4. render_and_save_image - render in Figma. To use this tool, you have to create unique image name for the rendered image with create_unique_image_name tool. 5. backup_plugin_folder - Backup plugin
Please implement the Figma plugin code following these steps:
1. Update the imageList in ui.html ("../figma-plugin-related/AdAgentBeta/ui.html") to include: {background_image_path}
2. Read the template from: "../figma-plugin-related/figma_plugin_template_code.js"
3. Generate the plugin code using the template and this layout specification: {layout output from textual director}
4. Save the generated code to code.js in: "../figma-plugin-related/AdAgentBeta/"
5. Render the image in Figma desktop and get the file path to the rendered image. Provide the file path output by the render_and_save_image tool. e.g. ../rendered/xxx.png
6. Copy the plugin folder "../figma-plugin-related/AdAgentBeta" under the backup folder "../figma-plugin-backup/" for backup purpose. Also copy the background image {full_background_image_path} under the backup plugin folder. The plugin folder’s backup name will be the background image name.
Follow the template structure and properly implement all text elements and CTA button with the specified styling and positioning.
Developer
You are iteratively evaluating the foreground elements (text, buttons, etc.) based on their fit with the background and your experience of a good banner ad in industry.
If the foreground elements are not perfect for produciton in industry, provide key FEEDBACK on ways to improve the foreground elements. If it looks good in overall, we should finish the development.
You should look at the current rendered image to spot any problems. The foreground strategy is intention while the rendered image is actually what has been implemented. So always look at the rendered image!
Here are some rules to write good key FEEDBACK:
  • If a feedback hasn’t been addressed for over two iterations, stop suggesting it again.
  • Carefully compare current rendered image with the foreground strategy to identify any abnormalities, such as foreground elements being partially outside the frame, misaligned, or distorted due to rendering issues. Write down any of these abnormalities in details. Carefully compare current rendered image with the background to examine foreground elements overlap with any salient areas of the background (e.g., key objects, patterns, colors). Write down any overlaps that might compromise readability or aesthetics in details. Carefully assess the placement of the logo image. Spot if the logo image is too small, too big, being distorted, or not placed in a prominent position. If need to resize the logo, suggest maintaining the aspect ratio. If the logo is hard to read (e.g. the color is similar to the background), suggest adding background to the logo image to make it more prominent. Carefully assess if the font type, font color, button color, button shape can be enhanced from a professional marketing expert’s perspective. Carefully assess if any overlaps among elements such as of text with the logo or button, or overlaps of button with the logo. If there are overlaps, suggest adjusting the position of the text, button, or logo to avoid overlaps. You can also propose better copywrite if you find it necessary. Use the provided history of previous iterations if it is not the first iteration. Avoid generating FEEDBACK identical with the FEEDBACK in previous rounds unless the issues raised before have not been solved. Ensure your FEEDBACK build on past refinements. Avoid any code blocks in your response. Use clear and concise human language. Spot the most critical issues first. Design Reviewer
  • Things to be careful are:
    • When you refine, avoid any potential overlapping of elements.
  • When you change the position of an element, consider step by step whether the relative positions of other elements—especially those that reference the changed element need to be adjusted to avoid disrupting the layout. When change the relative positions, you should always remember to update the position with the new relative positions, because relative position always affects positions. When changing elements, ensure the margins to the edges are considered. Check by calculating if the (position+size) exceeds ({background_width}, {background_height}) in horizontal and vertical dimension respectively.
  • Foreground Designer (Refinement)
    A.7 More Visual Results